# Understanding the Growth and Properties of Sputter-Deposited Phase-Change Superlattice Films

### v1

@@
Highly textured chalcogenide films have recently gained significant interest for phase-change memory applications. Several reports have highlighted that programming efficiency improves in devices featuring superlattice stacks, such as Ge2Sb2Te5/Sb2Te3. However, to be technologically relevant, these films must be deposited on foundry-scale wafers using processes compatible with back end of the line (BEOL) integration and complementary metal-oxide-semiconductor (CMOS) technology, such as, for example, sputter deposition. In this work, we present our observations on the influence of temperature, pressure, and seeding layer parameters on the sputter growth processes of superlattice films. By measuring various material properties, we construct a pseudo-phase diagram to illustrate the growth of both individual and superlattice films with different periodicities on technologically relevant substrates, namely SiO2 and carbon. These results provide important insights into the structure, intermixing and electro-optical properties of superlattice films, and identify optimal growth parameters critical for the large-scale manufacturability of the material.
@@ Results
Electrical characterization of superlattice films
The CSL films are expected to exhibit anisotropy between in-plane and cross-plane carrier transport properties[15]. To evaluate these, we measure the in-plane () and cross-plane () electrical resistivities using a modified-Transfer Length Measurements (m-TLM). Unlike standard TLM, which only estimate in-plane resistivity, m-TLM uniquely involves creating variable resistances by etching the film to different thicknesses (h). This is depicted in Figure 8a. In our m-TLM structures (see Supporting Information section I.4 for fabrication details), the CSL remains always capped and the contacts with electrodes are always intact. This is fundamentally different from previous studies[11], in which the films are exposed to ambience (thus, oxidation) during device fabrication. The total resistance measured between the two contacts can be written as
@@
Figure 8b shows an exemplar plot of as a function of the channel length (L) for a superlattice 96 nm thick progressively etched to vary the values of h. We firstly note that the increase in measured electrical resistance as a function of L is perfectly linear up to 40 m (the maximum pad distance), indicating excellent long-range uniformity of our films. is computed from the slopes of these curves, taking the value (average of extracted from each curve: 0.99 , 1.15 and 1.44 ). For , the intercepts () of these curves are plotted as a function of h and, since the intercept effectively equates to , can be extracted from the slope of the curve in Figure 8c, yielding . Therefore, as expected, a stark contrast between the cross-plane and in-plane resistivity, on the order of 100000x, is observed. This value, however, is notably higher than shown in the literature [11, 49, 24]. This is likely owing to our better controlled m-TLM structures and good film quality – the density of the vdW gaps.
@@ Conclusion
In summary, we have investigated the effects of growth parameters, such as pressure and temperature, on the sputter deposition of GST, Sb2Te3, and their superlattices (CSL). The seed layer of a specific thickness was critical for achieving oriented growth, with pressure being particularly important for Sb2Te3 due to its impact on Te desorption. For GST, deposition optimization enabled control over the stacking of the cubic phase without altering the target composition, with temperature being the most influential factor. A p,T diagram was developed to map optimal growth conditions for these films, resulting in highly textured and uniform growth of CSL films with varying periodicities on relevant substrates. Atomic characterization revealed local intermixing at the interfaces. Characterization of the optical refractive indices showed no dependency on periodicity. Electrical characterization using transfer length measurements demonstrated a significant contrast between cross-plane and in-plane resistivity. The quality of interfaces could be explained by deposition defined desorption kinetics, and films roughness. These results provide a directed methodology for wafer-scale growth of CSL.
@@ I.3 An understanding of the effect of pressure on CSL structure and morphology
I.4 Fabrication of TLM devices
In Section discussing electrical characterization in the main text, we presented data obtained from m-TLM to assess in-plane and cross-plane resistivity. Here, we provide a detailed description of the fabrication process.
The fabrication process for TLM test structures is schematized in Figure 13a. Initially, the substrate underwent the usual chemical cleaning procedure to prepare it for film deposition. Next, the CSL film was sputter-grown on the cleaned substrate. After allowing the film to cool down, it was capped with a 10 nm layer of tungsten (W) deposited at room temperature.
Following the film deposition, the chip was removed from the growth chamber for a first lithography step that, employing a negative resist to define the area of the TLM test structure. After the resist was developed, ion milling was performed to isolate the test structures as depicted in Figure 13b.
Subsequently, another lithography step with negative resist defined the pad area (100 m 100 m), leaving free of resist the contact area while the channel area remained covered. Subsequently, the chip was placed in a sputter tool for the deposition of100 nm W pads. Prior this deposition inverse sputter etch was conducted to remove any native W oxide that may have formed (1-2 nm).
The process then continued with a liftoff to define the final test structures. A profile of the result is reported in figure 13c. Ion milling concluded device fabrication etching the remaining W in the channel as well as a portion h of the superlattice. In this way, we fabricated 80 TLM test structures, 10 for each channel length (L=5, 10, 15, 20, 25, 30, 35, 40 m).
Here we report a table for comparison of electrical resistivity values:

### v2

@@
Highly textured chalcogenide films have recently gained significant interest for phase-change memory applications. Several reports have highlighted that programming efficiency improves in devices featuring superlattice stacks, such as Ge2Sb2Te5/Sb2Te3. However, to be technologically relevant, these films must be deposited on foundry-scale wafers using processes compatible with back end of the line (BEOL) integration and complementary metal-oxide-semiconductor (CMOS) technology, such as, for example, sputter deposition. In this work, we present our observations on the influence of temperature, pressure, and seeding layer parameters on the sputter growth processes of superlattice films. By measuring various material properties, we construct a pseudo-phase diagram to illustrate the growth of both individual and superlattice films with different periodicities on technologically relevant substrates, namely SiO2 and carbon. These results provide important insights into the structure, intermixing and electro-optical properties of superlattice films, and identify optimal growth parameters critical for the manufacturability via sputtering of the material.
@@ Results
To evaluate the electrical properties of CSL, we measured the in-plane () and cross-plane () electrical resistivity using a modified-Transfer Length Measurements (m-TLM). Unlike standard TLM, which only estimate in-plane resistivity, m-TLM uniquely involves creating variable resistances by etching the film to different thicknesses (h). This is depicted in Figure 8a. In our m-TLM structures (see Supporting Information section II for fabrication details), the CSL remains always capped and the contacts with electrodes are always intact. This is fundamentally different from previous studies[11], in which the film surface is exposed to ambiance (thus, oxidation) during device fabrication. The total resistance measured between the two contacts can be written as
@@
Figure 8b shows an exemplar plot of as a function of the channel length (L) for a superlattice 96 nm thick progressively etched to vary the values of h. We firstly note that the increase in measured electrical resistance as a function of L is perfectly linear up to 40 m (the maximum pad distance), indicating excellent long-range uniformity of our films. is computed from the slopes of these curves, taking the value (average of extracted from each curve: 0.99 , 1.15 and 1.44 ). This value is in line with superlattices data reported in literature, as shown in the supporting information.
In principle, modified TLM also allows to assess by plotting the intercepts () of these curves as a function of h. Calculating the with this approach in the supplementary information, we obtain, as expected, a value higher than . However, the measured is also significantly higher than values previously reported in the literature. This discrepancy could result from partial oxidation of the vertical pillars in the test structures, which may have affected the measurement. Further investigations are needed to confirm this, while a discussion and a comparison to literature are provided in the supplementary information.
@@ Conclusion
In summary, we have investigated the effects of growth parameters, such as pressure and temperature, on the sputter deposition of GST, Sb2Te3, and their superlattices (CSL). The seed layer of a specific thickness was critical for achieving oriented growth, with pressure being particularly important for Sb2Te3 due to its impact on Te desorption. For GST, deposition optimization enabled control over the stacking of the cubic phase, as well as the growth of t-GST, without altering the target composition, with temperature being the most influential factor. This is particularly relevant in light of renewed interest in textured GST as active material[42, 49]. A p,T diagram was developed to map optimal growth conditions for these films, resulting in highly textured and uniform growth of CSL films with varying periodicities on relevant substrates. CSL growth was also achieved using a seed layer deposited at high T, useful to minimize potential delamination. Atomic characterization revealed local intermixing at the interfaces. Characterization of the optical refractive indices showed no dependency on periodicity. The quality of interfaces could be explained by deposition defined desorption kinetics and film roughness. These results provide a directed methodology for sputter growth of CSL as well as textured Sb2Te3 and GST.
@@ I.3 An understanding of the effect of pressure on CSL structure and morphology
II Fabrication of TLM devices and cross plane resistivity
In Section discussing electrical characterization in the main text, we presented data obtained from m-TLM to assess in-plane and cross-plane resistivity. Here, we provide a detailed description of the 8 fabrication steps schematized in Figure 13. Initially, the substrate underwent the usual chemical cleaning procedure to prepare it for film deposition. Next, the CSL film was sputter-grown on the cleaned substrate. After allowing the film to cool down, it was capped with a 10 nm layer of tungsten (W) deposited at room temperature (1). This W layer will be the starting point for the contacts and, since it is deposited in situ, it effectively prevents CSL surface exposure to atmosphere (thus oxidation).
Following film deposition, the chip was removed from the growth chamber for a first lithography step that employed a negative resist to define the area of the TLM test structure (2). After the resist was developed, an ion milling etching was performed to isolate the test structures (3). The resist was then removed (4).
Subsequently, another lithography step (5) with negative resist was carried out to define the pad area (100 m 100 m). This consisted of leaving the contact area free from resist while the channel area remained covered. Subsequently, the chip was placed in a sputter tool for the deposition of100 nm W pads (6). Prior this deposition inverse sputter etch was conducted to remove any native W oxide that may have formed (1-2 nm).
The process then continued with a liftoff that removed the resists and defined the final test structures (7). A final ion milling step (8) ended the fabrication of the device, etching the remaining W in the channel as well as a portion h of the superlattice. In this way, we fabricated 80 TLM test structures, 10 for each channel length (L=5, 10, 15, 20, 25, 30, 35, 40 m).
For , the intercepts () of the curves are plotted as a function of h and, since the intercept effectively equates to . Consequently, can be extracted from the slope of the curve in Figure 14, yielding . This translates in a stark contrast between the cross-plane and in-plane resistivity, on the order of 100000 x. However, as summarized in Table 14, the / ratio observed here is significantly higher than previous reports in the literature [11, 53, 24], primarily due to an unusually high cross-plane resistivity. While a substantial difference between in-plane and cross-plane conductivity is expected, such an elevated value could be influenced by additional factors. For instance, all the sidewalls of the vertical pillars, which contribute to the cross-plane resistivity, are directly exposed to the atmosphere, increasing the surface area susceptible to oxidation. This could compromise the accuracy of the cross-plane resistivity measurement. Therefore, to validate the reported value, further investigations are necessary to rule out the influence of such effects. Conversely, the in plane counterpart is less affected as the portion of film involved in the evaluation of exposes much less area to the ambiance. As a result, the measured value is in line with previous reports (see table 14).